OpenStudy (anonymous):

Did I do this Algebra problem correctly? Write an equation of the line that passes through each point with the given slope. (-5, 5) m=-2 I think the answer is y=-2x+5 right

OpenStudy (anonymous):

Last time I asked this, I received two different responses

OpenStudy (anonymous):

is it y=-2x+5 or y=-2x-5?

OpenStudy (mertsj):

Use point slope form: \[y-y _{1}=m(x-x _{1})\] Replace \[y _{1}\] with 5. Replace m with -2. Replace \[x _{1}\] with -5 . Simplify. That will be your equation.

OpenStudy (anonymous):

okay, so it's positive 5. thank you

OpenStudy (mertsj):

yw

OpenStudy (anonymous):

can you help me with another thing?

OpenStudy (anonymous):

write in standard form y+6=(x+4) the answer is X+Y=-2, correct? PLEASE just say yes or no.

OpenStudy (mertsj):

First I want to make sure you did this one correctly. Did you distribute the -2 correctly?

OpenStudy (mertsj):

no

OpenStudy (anonymous):

Alright, thanks. That's a huge help.

OpenStudy (anonymous):

how do i do it, then? :)

OpenStudy (mertsj):